# Higher-Dimensional Moving Averages and Submanifold Genericity

### v1

@@ 1.4. Acknowledgments
The authors are indebted to Alexander Gorodnik for suggesting to read [BJR90], to Dmitry Kleinbock for pointing out Remark 1.11, and to Amos Nevo for many fruitful conversations which eventually lead to this paper.
2. Proof of Theorems 1.1 and 1.7 and of Corollaries 1.3 and 1.8
@@
First, we observe that, if a box $B_{k}$ contains both integer vectors whose $i$-th component is negative and integer vectors whose $i$-th component is positive for some $k$ and $i$, then it must be $n_{ki}<0$ and $l_{ki}>|n_{ki}|$. In this case, we may write
@@
If $\mathcal{B}^{\prime}$ is the collection of boxes where $B_{k}$ is replaced by $B_{k}^{i-}$ and $B_{k}^{i+}$, it is obvious that ${M_{\mathcal{B}}}f\leq M_{\mathcal{B}^{\prime}}f$ for any $f\in\textup{L}^{p}(X)$. Moreover, for any $k$ and $i$ such that $l_{ki}>|n_{ki}|$ and any $\lambda\in\mathbb{N}$ we have that
#$\#\{x\in\mathbb{Z}:|x-n_{ki}|\leq\alpha(\lambda-l_{ki})\}\leq(1+\alpha)\lambda,$
so that the boxes $B_{k}$ that are broken in two or more parts do not contribute to the validity of Condition ($C_{i}$). Hence, by working in each orthant separately and by replacing $T_{i}$ with $T_{i}^{-1}$ if necessary, we may always assume that $n_{ki}\geq 0$ for all $k$ and $i$.
Take $f\in\textup{L}^{p}(X)$ for fixed $p>1$ and observe that
${M_{\mathcal{B}}}f(x)$ $=\sup_{B_{k}\in\mathcal{B}}\frac{1}{l_{k1}\dotsm l_{kd}}\sum_{\boldsymbol{j}\in B_{k}\cap\mathbb{Z}^{d}}\left|f\left(T_{1}^{j_{1}}\dotsb T_{d}^{j_{d}}x\right)\right|$
$\leq\sup_{(n_{k1},l_{k1})\in\Omega_{1}}\dotsb\sup_{(n_{kd},l_{kd})\in\Omega_{d}}\frac{1}{l_{k1}\dotsm l_{kd}}\sum_{j_{1}=0}^{l_{k1}-1}\dotsb\sum_{j_{d}=0}^{l_{kd}-1}\left|f\left(T_{1}^{n_{k1}+j_{1}}\dotsb T_{d}^{n_{kd}+j_{d}}x\right)\right|$
$\leq\sup_{(n_{k1},l_{k1})\in\Omega_{1}}\frac{1}{l_{k1}}\sum_{j_{1}=0}^{l_{k1}-1}\left|\ \dotsb\sup_{(n_{kd},l_{kd})\in\Omega_{d}}\frac{1}{l_{kd}}\sum_{j_{d}=0}^{l_{kd}-1}\left|f\left(T_{1}^{n_{k1}+j_{1}}\dotsb T_{d}^{n_{kd}+j_{d}}x\right)\right|\right|$
@@
$M_{\Omega_{i}}f(x):=\sup_{(n_{ki},l_{ki})\in\Omega_{i}}\frac{1}{l_{ki}}\sum_{j=0}^{l_{ki}-1}\left|f\left(T^{n_{ki}+j}x\right)\right|$
@@
By [BJR90, Theorem 1], each operator $M_{\Omega_{i}}$ is of strong type $(p,p)$. Then ${M_{\mathcal{B}}}$ is also of strong type $(p,p)$.
2.2. Proof of Theorem 1.7
The proof in the case $d>1$ relies once again on the case $d=1$ and is analogous to that of Theorem 1.1. We therefore leave it to the reader.
In the case $d=1$ we modify the proof of [BJR90, Theorem 1]. This requires some additional work. In what follows, for any $p\geq 1$ and any measure space $Y$ we denote by $\mathscr{L}^{p}(Y)$ the set of $p$-integrable functions from $Y$ to $\mathbb{R}$.
For any $f\in\mathscr{L}^{1}(X)$, $T>0$, $x\in X$, and $t>0$ put
@@
Then $f_{T,x}\in\mathscr{L}^{1}(\mathbb{R})$. Let $f_{T,x}^{*}$ denote the one-sided Hardy-Littlewood maximal function associated to $f_{T,x}$, that is
@@
where the bounds of the integral are to be inverted if $r<0$. By the Hardy-Littlewood maximal inequality [Tao11, Lemma 1.6.16] we have that for any $\lambda>0$
@@
Let us now fix $\lambda>0$ and a function $f\in\mathscr{L}^{1}(X)$. Then
@@
Finally, for any $x\in X$ let
$Y_{\varepsilon}(x):=\{\xi:(x,\xi)\in Y_{\varepsilon}\}.$
Then we have the following.
@@
(2.3) $[t+w_{k},t+w_{k}+s_{k})\subset\left\{\xi:f_{T_{\varepsilon},x}^{*}(\xi)>\lambda\right\}.$
@@
where $A$ and $\alpha$ are the constants in Condition ($\tilde{C}_{i}$) (for $i=1$) and $C_{\alpha}$ is an absolute constant only depending on $\alpha$.
@@
By the Vitali Covering Lemma (see [Coh13, Theorem 6.2.1]), there exists a countable sub-collection of disjoint open intervals $(\xi_{i},\xi_{i}+r_{i})$ (where $r_{i}=r_{\xi_{i}}$ for brevity) such that
@@
then, there must be $i$ such that $t+w_{k}+s_{k}/2\in(\xi_{i}-2r_{i},\xi_{i}+3r_{i})$. Hence, $63r_{i}\geq s_{k}/$ 6. In any case, we have that there exists $i$ such that $r_{i}\geq s_{k}/18$ and $(\xi_{i}-2r_{i},\xi_{i}+3r_{i})\cap[t+w_{k},t+w_{k}+s_{k})\neq\emptyset$. For this $i$ we therefore have that
@@
On combining Lemma 2.2 and (2.1), we find that for any $x\in X$ it holds that
@@
On letting $\varepsilon\to 0$, we conclude that ${N_{\mathcal{B}}}$ is of weak type $(1,1)$. Since ${N_{\mathcal{B}}}$ is bounded from $\mathrm{L}^{\infty}$ to $\mathrm{L}^{\infty}$, by the Marcinkiewicz Interpolation Theorem [Gra14, Theorem 1.3.2], ${N_{\mathcal{B}}}$ is of strong type $(p,p)$ for $p>1$.
2.3. Proof of Corollaries 1.3 and 1.8
The strategy to prove an ergodic theorem given a maximal inequality is standard (see for example [EW11, Section 2.6.5]). We therefore only give a sketch of proof for the discrete case. Let us start with an observation.
@@
If for some $1\leq i\leq d$ Condition ($C_{i}$) is satisfied, then it must be $l_{ki}\to\infty$. In fact, if there is a sequence $k_{r}$ such that $l_{k_{r}i}\leq C$ for a given constant $C\geq 1$, then for all integer $\lambda\geq C$ and all $r$ we have that
$(n_{k_{r}i},\lambda)\in\Omega_{i}^{(\alpha)}(\lambda),$
whence #$\#\Omega_{i}^{(\alpha)}(\lambda)=\infty$.
@@
Let $f\in\textup{L}^{2}(X)$. Then there exists a function $f^{\prime}\in\textup{L}^{2}(X)$ that is invariant under the transformations $T_{1},\dotsc,T_{d}$, such that
$\|A_{k}f-f^{\prime}\|_{2}\to 0.$
@@
For $i=1,\dotsc,d$ define $U_{T_{i}}f:=f\circ T_{i}$ and let
$I:=\{g\in\textup{L}^{2}(X):U_{T_{i}}g=g\mbox{ for }i=1,\dotsc,d\}.$
@@
$B_{i}:=\{U_{T_{i}}g-g:g\in\textup{L}^{2}(X)\}.$
Since $l_{ki}\to\infty$ for $i=1,\dotsc,d$ (see Remark 2.3) it is clear that for any $h\in B_{i}$ we have that $\|A_{k}h\|_{2}\to 0$ as $k\to\infty$. Then the conclusion follows as in [EW11, Theorem 2.21]. ∎
Let $f\in\textup{L}^{\infty}(X)$. By Lemma 2.4, we know that $A_{k}f$ converges to an invariant function $f^{\prime}$ in $\textup{L}^{2}(X)$. Now, for every measurable $B\subset X$ we have that
$\langle A_{k}f,\chi_{B}\rangle\leq\|f\|_{\infty}\cdot\mu(B).$
Hence, the same must be true if $A_{k}f$ is replaced by $f^{\prime}$. This shows that $f^{\prime}\in\textup{L}^{\infty}(X)$.
@@
$A_{k}\circ A_{h}f=A_{k}f+O_{\boldsymbol{n}_{h},\boldsymbol{l}_{h}}\left((l_{k1}\dotsm l_{kd})^{-1}\|f\|_{\infty}\right).$
@@
We have that
$A_{k}\circ A_{h}f$ $=\frac{1}{l_{k1}\dotsm l_{kd}}\sum_{j_{1}^{\prime}=0}^{l_{k1}-1}\dotsb\sum_{j_{d}^{\prime}=0}^{l_{kd}-1}$
$\hskip 56.9055pt\frac{1}{l_{h1}\dotsm l_{hd}}\sum_{j_{1}=0}^{l_{h1}-1}\dotsb\sum_{j_{d}=0}^{l_{hd}-1}f\left(T_{1}^{n_{h1}+n_{k1}+j_{1}+j_{1}^{\prime}}\dotsb T_{d}^{n_{hd}+n_{kd}+j_{d}+j_{d}^{\prime}}x\right)$
$=\frac{1}{l_{h1}\dotsm l_{hd}}\sum_{j_{1}=0}^{l_{h1}-1}\dotsb\sum_{j_{d}=0}^{l_{hd}-1}$
$\hskip 56.9055pt\frac{1}{l_{k1}\dotsm l_{kd}}\sum_{j_{1}^{\prime}=n_{h1}+j_{1}}^{l_{k1}-1+n_{h1}+j_{1}}\dotsb\sum_{j_{d}^{\prime}=n_{hd}+j_{d}}^{l_{kd}-1+n_{hd}+j_{d}}f\left(T_{1}^{n_{k1}+j_{1}^{\prime}}\dotsb T_{d}^{n_{kd}+j_{d}^{\prime}}x\right)$
$=A_{k}f+O_{\boldsymbol{n}_{h},\boldsymbol{l}_{h}}\left((l_{k1}\dotsm l_{kd})^{-1}\|f\|_{\infty}\right),$
@@
Let us show point-wise convergence in $\textup{L}^{\infty}(X)$. Let $f\in\textup{L}^{\infty}(X)$ and let $f^{\prime}$ be the function found in Lemma 2.4. Note that for any $k$ it holds that $A_{k}f^{\prime}=f^{\prime}$, since $f^{\prime}$ is invariant. Fix $\varepsilon,\delta>0$ and pick $h$ so large that $\|A_{h}f-f^{\prime}\|_{2}\leq\delta$. Then, by Lemma 2.5, Remark 2.3, and Theorem 1.7, we have that
$\mu\left(x:\limsup_{k}|A_{k}f-f^{\prime}|>\varepsilon\right)=\mu\left(x:\limsup_{k}|A_{k}\circ A_{h}f-A_{k}f^{\prime}|>\varepsilon\right)\\ \leq\mu\left(x:{M_{\mathcal{B}}}(A_{h}f-f^{\prime})>\varepsilon\right)\ll\|A_{h}f-f^{\prime}\|_{2}\leq\delta.$
@@
3. Proof of Theorems 1.5 and 1.9
In this section we show that, provided Condition ($!C_{i}$) (respectively ($!\tilde{C}_{i}$)) holds for some $1\leq i\leq d$, the operators $A_{k}$ (respectively $R_{k}$) enjoy the "strong sweeping out" property. This is a direct application of [BJR90, Theorem 3], which we recall below for the convenience of the reader.
@@
When $T_{k}=A_{k}$ (or $T_{k}=R_{k}$), the first two assumptions in Theorem 3.1 are trivially true, while the third one follows by our hypothesis on the transformations $T_{1},\dotsc,T_{d}$ (or the flow $U_{\boldsymbol{t}}$). Hence, it is enough for us to verify that ($\star$3.1) holds. Note also that, when $f$ is the characteristic function of a set in $X$, $M_{n}f$ coincides with $M_{\mathcal{B}^{\prime}}f$ (or $N_{\mathcal{B}^{\prime}}f$), where =$\mathcal{B}^{\prime}=\{B_{k}\in\mathcal{B}:k\geq n\}$.
3.1. Proof of Theorem 1.5
Take $\alpha=1$ and assume, without loss of generality, that ($!C_{i}$) holds for $i=1$. Fix $p\in\mathbb{N}$ and choose an integer $\lambda_{p}$ so that #$\#\Omega_{1}^{(1)}(\lambda_{p})\geq p\cdot(4\lambda_{p}+1)$. Recall that, by definition of $\Omega_{1}^{(1)}(\lambda_{p})$, for any $z\in\Omega_{1}^{(1)}(\lambda_{p})$ there exists $(n_{k1},l_{k1})\in\Omega_{1}$ such that
(3.1) $|z-n_{k1}|\leq\lambda_{p}-l_{k1}\leq\lambda_{p}.$
In view of this, we may write
$\Omega_{1}^{(1)}(\lambda_{p})=\bigcup_{k}C_{k}^{(1)}(\lambda_{p}),$
where
$C_{k}^{(1)}(\lambda_{p}):=\{x\in\mathbb{Z}:|x-n_{k1}|\leq(\lambda_{p}-l_{k1})\}.$
Choose $K_{p}\geq 1$ large enough so that
#$\#\left(\bigcup_{k\leq K_{p}}C_{k}^{(1)}(\lambda_{p})\right)\geq p\cdot(4\lambda_{p}+1)$
and put
$\Delta:=\bigcup_{k\leq K_{p}}C_{k}^{(1)}(\lambda_{p}).$
Then for each $z\in\Delta$ there exists $k\leq K_{p}$ such that (3.1) holds. Note that it must be $\lambda_{p}\geq l_{k1}$, so that for all $j=0,\dotsc,l_{k1}-1$ we have that
(3.2) $|-z+n_{k1}+j|\leq\lambda_{p}+\lambda_{p}=2\lambda_{p}.$
Now, let
$N_{1}:=2\lambda_{p}+\sup\Delta$
and
$N_{j}:=\sup_{k\leq K_{p}}|n_{kj}+l_{kj}|$
for $j=2,\dotsc,d$. Form a Rokhlin tower as in [KW72, Theorem 1] with parameters $N_{1},3N_{2},\dotsc,3N_{d}$. Then
$\mu\left(\bigcup_{j_{1}=0}^{N_{1}-1}\bigcup^{3N_{2}-1}_{j_{2}=0}\ldots\bigcup_{j_{d}=0}^{3N_{d}-1}T_{1}^{j_{1}}T_{2}^{j_{2}}\ldots T_{d}^{j_{d}}(B)\right)>1-\delta$
for some positive-measure set $B\subset X$ and some small $\delta>0$, where the union is disjoint. Define
$H_{p}:=\bigcup_{j_{1}=N_{1}-4\lambda_{p}-1}^{N_{1}-1}\bigcup^{3N_{2}-1}_{j_{2}=0}\dotsb\bigcup_{j_{d}=0}^{3N_{d}-1}T_{1}^{j_{1}}T_{2}^{j_{2}}\ldots T_{d}^{j_{d}}(B).$
Fix $z\in\Delta$ and
$x\in F_{p}:=\bigcup^{2N_{2}-1}_{j_{2}=N_{2}}\dotsb\bigcup_{j_{d}=N_{d}}^{2N_{d}-1}T_{1}^{N_{1}-2\lambda_{p}-1}T_{2}^{j_{2}}\ldots T_{d}^{j_{d}}(B).$
Then, by (3.2), for some $k\leq K_{p}$ it holds that
$N_{1}-4\lambda_{p}-1\leq(N_{1}-2\lambda_{p}-1)-z+n_{k1}+j\leq N_{1}-1$
for all $j=0,\dotsc,l_{k1}-1$. Since $|n_{kj}+l_{kj}|\leq N_{j}$ for $j=2,\dotsc,d$, we conclude that
$M_{\mathcal{B}}\chi_{H_{p}}(T_{1}^{-z}x)\geq\frac{1}{l_{k1}\dotsm l_{kd}}\sum_{j_{1}=0}^{l_{k1}-1}\dotsb\sum_{j_{d}=0}^{l_{kd}-1}\chi_{H_{p}}\left(T_{1}^{-z+n_{k1}+j_{1}}T_{2}^{n_{k2}+j_{2}}\dotsb T_{d}^{n_{kd}+j_{d}}x\right)=1.$
This implies that
$\left\{{M_{\mathcal{B}}}\chi_{H_{p}}>1-\varepsilon\right\}\supset\bigcup_{z\in\Delta}T_{1}^{-z}F_{p},$
whence
$\mu\left(M_{\mathcal{B}}\chi_{H_{p}}>1-\varepsilon\right)\geq\#\Delta\cdot N_{2}\ldots N_{d}\cdot\mu(B).$
Moreover, it is clear that
$\mu(H_{p})=3^{d-1}(4\lambda_{p}+1)\cdot N_{2}\ldots N_{d}\cdot\mu(B).$
Combining these two observations, we find that
$\frac{\mu\left(M_{\mathcal{B}}\chi_{H_{p}}>1-\varepsilon\right)}{\mu(H_{p})}\geq\frac{\#\Delta\cdot N_{2}\ldots N_{d}\cdot\mu(B)}{3^{d-1}(4\lambda_{p}+1)\cdot N_{2}\ldots N_{d}\cdot\mu(B)}\geq\frac{p}{3^{d-1}},$
showing that ($\star$3.1) holds for the operator $M_{\mathcal{B}}$. To conclude, observe that if :=$\mathcal{B}^{\prime}:=\{B_{k}\in\mathcal{B}:k\geq n\}$, then ($!C_{i}$) holds for the collection $\mathcal{B}^{\prime}$ and the above argument also applies to the operator $M_{\mathcal{B}^{\prime}}$.
3.2. Proof of Theorem 1.9
The continuous case follows from a similar argument, based on a Rokhlin tower construction for aperiodic flows proved in [Lin75], which we recall below.
@@
[Lin75, Theorem 1] Let $U_{\boldsymbol{t}}=U_{t_{1},\ldots,t_{d}}$ be a $d$-dimensional measure-preserving aperiodic flow on $(X,\mu)$. Let $L_{1},\ldots,L_{d},\delta>0$ and let $Q=Q_{L}:=[0,L_{1})\times\ldots\times[0,L_{d})$. Then there exists a set $B\subset X$ with the following properties:

### v2

@@ 1.3. Acknowledgments
The authors are indebted to Alexander Gorodnik for suggesting to read [BJR90], to Dmitry Kleinbock for pointing out Remark 1.8, and to Amos Nevo for many fruitful conversations which eventually led to this paper.
2. Proof of Theorem 1.2 and of Corollary 1.3
@@
First, we observe that, if a box $B_{k}$ contains both vectors whose $i$-th component is negative and vectors whose $i$-th component is positive for some $k$ and $i$, then it must be $w_{ki}<0$ and $s_{ki}>|w_{ki}|$. In this case, we may write
@@
If $\mathcal{B}^{\prime}$ is the collection of boxes where $B_{k}$ is replaced by $B_{k}^{i-}$ and $B_{k}^{i+}$, it is obvious that ${N_{\mathcal{B}}}f\leq N_{\mathcal{B}^{\prime}}f$ for any $f\in\textup{L}^{p}(X)$. Moreover, for any $k$ and $i$ such that $s_{ki}>|w_{ki}|$ and any $\lambda>0$ we have that
$\mathrm{Leb}(\{x\in\mathbb{R}:|x-w_{ki}|\leq\alpha(\lambda-s_{ki})\})\leq(1+\alpha)\lambda,$
so that the boxes $B_{k}$ that are broken in two or more parts do not contribute to the validity of Condition ($\tilde{C}_{i}$). Hence, by working in each orthant separately and by replacing $U_{t_{i}}$ with $U_{t_{i}}^{-1}$ if necessary, we may always assume that $w_{ki}\geq 0$ for all $k$ and $i$.
Now, the proof for $d>1$ relies on the case $d=1$ (with $s_{k}>0$). To see this let $U_{t}^{(i)}:=U_{t\boldsymbol{e}_{i}}$, where $\boldsymbol{e}_{i}$ denotes the $i$-th vector of the standard basis in $\mathbb{R}^{d}$. Take $f\in\textup{L}^{p}(X)$ for fixed $p>1$ and observe that
${N_{\mathcal{B}}}f(x)$ $=\sup_{B_{k}\in\mathcal{B}}\frac{1}{s_{k1}\dotsm s_{kd}}\int_{B_{k}}\left|f\left(U^{(1)}_{t_{1}}\dotsb U^{(d)}_{t_{d}}x\right)\right|\,\textup{d}\boldsymbol{t}$
$\leq\sup_{(w_{k1},s_{k1})\in\Omega_{1}}\dotsb\sup_{(w_{kd},s_{kd})\in\Omega_{d}}\frac{1}{s_{k1}\dotsm s_{kd}}\int_{t_{1}=w_{k1}}^{w_{k1}+s_{k1}}\dotsb\int_{t_{d}=w_{kd}}^{w_{kd}+s_{kd}}\left|f\left(U^{(1)}_{t_{1}}\dotsb U^{(d)}_{t_{d}}x\right)\right|\,\textup{d}\boldsymbol{t}$
$\leq\sup_{(w_{k1},s_{k1})\in\Omega_{1}}\frac{1}{s_{k1}}\int_{t_{1}=w_{k1}}^{w_{k1}+s_{k1}}\left|\ \dotsb\sup_{(w_{kd},s_{kd})\in\Omega_{d}}\frac{1}{s_{kd}}\int_{t_{d}=w_{kd}}^{w_{kd}+s_{kd}}\left|f\left(U^{(1)}_{t_{1}}\dotsb U^{(d)}_{t_{d}}x\right)\right|\right|\,\textup{d}\boldsymbol{t}$
@@
$M_{\Omega_{i}}f(x):=\sup_{(w_{ki},s_{ki})\in\Omega_{i}}\frac{1}{s_{ki}}\int_{t=w_{ki}}^{w_{ki}+s_{ki}}\left|f\left(U^{(i)}_{t}x\right)\right|\,\textup{d}t$
@@
Now, by the case $d=1$, each operator $M_{\Omega_{i}}$ is of strong type $(p,p)$. Then $M_{\mathcal{B}}$} is also of strong type $(p,p)$.
For the case $d=1$ we modify the proof of [BJR90, Theorem 1]. We stress once again that the conclusion for $d=1$ can also be deduced by the arguments in [Ferrando_1995], however, modifying [BJR90, Theorem 1] gives a much simpler proof, which we present to keep the paper self-contained.
Recall that a family of intervals $\mathcal{B}=\{[w_{k},w_{k}+s_{k})\}$ with $w_{k}\geq 0$ and $s_{k}>0$ is given and that this family satisfies Condition ($\tilde{C}_{i}$). Our goal is to show that the operator
$N_{\mathcal{B}}f(x):=\sup_{k}\frac{1}{s_{k}}\int_{0}^{s_{k}}f(U_{w_{k}+t}x)\,\textup{d}t$
is of strong type $(p,p)$ for any $p>1$. We will in fact show that it is of weak type $(1,1)$ and conclude by a classical interpolation argument.
In what follows, for any $p\geq 1$ and any measure space $Y$ we denote by $\mathscr{L}^{p}(Y)$ the set of $p$-integrable functions from $Y$ to $\mathbb{R}$. For any $f\in\mathscr{L}^{1}(X)$, $T>0$, $x\in X$, and $t>0$ we also put
@@
Then, clearly, $f_{T,x}\in\mathscr{L}^{1}(\mathbb{R})$. Denote by $f_{T,x}^{*}$ the one-sided Hardy-Littlewood maximal function associated to $f_{T,x}$, that is
@@
where the bounds of the integral are to be inverted if $r<0$. Then, by the Hardy-Littlewood maximal inequality [Tao11, Lemma 1.6.16], we have that for any $\lambda>0$
@@
Let us now fix $\lambda>0$ and a function $f\in\mathscr{L}^{1}(X)$. Recall that we aim to show that
(2.2) $\mu\left(x:{N_{\mathcal{B}}}f(x)>\lambda\right)\leq\frac{\textup{const.}}{\lambda}\|f\|_{1}.$
First, observe that
@@
Our goal will now be to bound from above the measure of the set $Y_{\varepsilon}$. To do so, we consider sections
$Y_{\varepsilon}(x):=\{\xi:(x,\xi)\in Y_{\varepsilon}\}$
for $x\in X$. In the next two lemmas, we obtain a bound for $\textup{Leb}(Y_{\varepsilon}(x))$ in terms of the quantity
$\textup{Leb}\left(\xi:f_{T,x}^{*}(\xi)>\lambda\right)$
for fixed $x$. Thanks to (2.1) and (2.3), this will easily lead to (2.2) (details explained in the sequel).
@@
(2.4) $t+w_{k},t+w_{k}+s_{k})\subset\left\{\xi:f_{T_{\varepsilon},x}^{*}(\xi)>\lambda\right\}.$
@@
where AA and $\alpha$ are the constants in Condition ($\tilde{C}_{i}$) (for $i=1$) and $C_{\alpha}$ is an absolute constant only depending on $\alpha$.
@@
By the Vitali Covering Lemma (see [Cohn13, Theorem 6.2.1]), there exists a countable sub-collection of disjoint open intervals $(\xi_{i},\xi_{i}+r_{i})$ (where $r_{i}=r_{\xi_{i}}$ for brevity) such that
@@
then, there must be $i$ such that $t+w_{k}+s_{k}/2\in(\xi_{i}-2r_{i},\xi_{i}+3r_{i})$. Hence, $3r_{i}\geq s_{k}/6$. In any case, we have that there exists $i$ such that $r_{i}\geq s_{k}/18$ and $(\xi_{i}-2r_{i},\xi_{i}+3r_{i})\cap[t+w_{k},t+w_{k}+s_{k})\neq\emptyset$. For this $i$ we therefore have that
@@
On combining (2.1) and Lemma 2.2, we find that for any $x\in X$ it holds that
@@
On letting $\varepsilon\to 0$, we conclude that $N_{\mathcal{B}}$} is of weak type $(1,1)$. Since $N_{\mathcal{B}}$} is bounded from $\mathrm{L}^{\infty}$ to $\mathrm{L}^{\infty}$, by the Marcinkiewicz Interpolation Theorem [Grafakos, Theorem 1.3.2], $N_{\mathcal{B}}$} is of strong type $(p,p)$ for $p>1$.
2.2. Proof of Corollary 1.3
The strategy to prove an ergodic theorem given a maximal inequality is standard (see for example [EW11, Section 2.6.5]). We therefore only sketch the proof. Let us start with an observation.
@@
If for some $1\leq i\leq d$ Condition ($\tilde{C}_{i}$) is satisfied, then it must be $s_{ki}\to\infty$. In fact, if there is a sequence $k_{r}$ such that $s_{k_{r}i}\leq C$ for a given constant $C\geq 1$, then for all integer $\lambda\geq C$ and all $r$ we have that
$\{(w_{k_{r}i}+x,\lambda):-\alpha\cdot(\lambda-C)\leq x\leq\alpha\cdot(\lambda-C)\}\subset\Omega_{i}^{(\alpha)}(\lambda),$
whence $\mathrm{Leb}\left(\Omega_{i}^{(\alpha)}(\lambda)\right)=\infty$.
@@
Let $f\in\textup{L}^{2}(X)$. Then there exists a function $f^{\prime}\in\textup{L}^{2}(X)$ that is invariant under the flow $U_{\boldsymbol{t}}$, such that
$\|R_{k}f-f^{\prime}\|_{2}\to 0.$
@@
For $i=1,\dotsc,d$ define $\mathcal{U}_{U_{t_{i}}}f:=f\circ U_{t_{i}}$ and let
$I:=\{g\in\textup{L}^{2}(X):\mathcal{U}_{U_{t_{i}}}g=g\mbox{ for }i=1,\dotsc,d\}.$
@@
$B_{i}:=\{\mathcal{U}_{U_{t_{i}}}g-g:g\in\textup{L}^{2}(X)\}.$
Since $s_{ki}\to\infty$ for $i=1,\dotsc,d$ (see Remark 2.3) it is clear that for any $h\in B_{i}$ we have that $\|R_{k}h\|_{2}\to 0$ as $k\to\infty$. Then the conclusion follows as in [EW11, Theorem 2.21]. ∎
Let $f\in\textup{L}^{\infty}(X)$. By Lemma 2.4, we know that $R_{k}f$ converges to an invariant function $f^{\prime}$ in $\textup{L}^{2}(X)$. Now, for every measurable $B\subset X$ we have that
$\langle R_{k}f,\chi_{B}\rangle\leq\|f\|_{\infty}\cdot\mu(B).$
Hence, the same must be true if $R_{k}f$ is replaced by $f^{\prime}$. This shows that $f^{\prime}\in\textup{L}^{\infty}(X)$.
@@
$R_{k}\circ R_{h}f=R_{k}f+O_{\boldsymbol{w}_{h},\boldsymbol{s}_{h}}\left((s_{k1}\dotsm s_{kd})^{-1}\|f\|_{\infty}\right).$
@@
In this proof, we once again denote by $U_{t}^{(i)}$ the one-dimensional flow $U_{t\boldsymbol{e}_{i}}$ for $i=1,\dotsc,d$. We have that
$R_{k}\circ R_{h}f$ $=\frac{1}{s_{k1}\dotsm s_{kd}}\int_{t_{1}^{\prime}=0}^{t_{1}^{\prime}=s_{k1}}\dotsb\int_{t_{d}^{\prime}=0}^{t_{d}^{\prime}=s_{kd}}$
$\hskip 14.22636pt\frac{1}{s_{h1}\dotsm s_{hd}}\int_{t_{1}=0}^{t_{1}=s_{h1}}\dotsb\int_{t_{d}=0}^{t_{d}=s_{hd}}f\left(U^{(1)}_{w_{h1}+w_{k1}+t_{1}+t_{1}^{\prime}}\dotsb U^{(d)}_{w_{hd}+w_{kd}+t_{d}+t_{d}^{\prime}}x\right)\,\textup{d}\boldsymbol{t}\,\textup{d}\boldsymbol{t}^{\prime}$
$=\frac{1}{s_{h1}\dotsm s_{hd}}\int_{t_{1}=0}^{t_{1}=s_{h1}}\dotsb\int_{t_{d}=0}^{t_{d}=s_{hd}}$
$\hskip 14.22636pt\frac{1}{s_{k1}\dotsm s_{kd}}\int_{t_{1}^{\prime}=w_{h1}+t_{1}}^{t^{\prime}_{1}=s_{k1}+w_{h1}+t_{1}}\dotsb\int_{t_{d}^{\prime}=w_{hd}+t_{d}}^{t_{d}^{\prime}=s_{kd}+w_{hd}+t_{d}}f\left(U^{(1)}_{w_{k1}+t_{1}^{\prime}}\dotsb U^{(d)}_{w_{kd}+t_{d}^{\prime}}x\right)\,\textup{d}\boldsymbol{t}^{\prime}\,\textup{d}\boldsymbol{t}$
$=R_{k}f+O_{\boldsymbol{w}_{h},\boldsymbol{s}_{h}}\left((s_{k1}\dotsm s_{kd})^{-1}\|f\|_{\infty}\right),$
@@
Let us show point-wise convergence in $\textup{L}^{\infty}(X)$. Let $f\in\textup{L}^{\infty}(X)$ and let $f^{\prime}$ be the function found in Lemma 2.4. Note that for any $k$ it holds that $R_{k}f^{\prime}=f^{\prime}$, since $f^{\prime}$ is invariant. Fix $\varepsilon,\delta>0$ and pick $h$ so large that $\|R_{h}f-f^{\prime}\|_{2}\leq\delta$. Then, by Lemma 2.5, Remark 2.3, and Theorem 1.2, we have that
$\mu\left(x:\limsup_{k}|R_{k}f-f^{\prime}|>\varepsilon\right)=\mu\left(x:\limsup_{k}|R_{k}\circ R_{h}f-R_{k}f^{\prime}|>\varepsilon\right)\\ \leq\mu\left(x:{N_{\mathcal{B}}}(R_{h}f-f^{\prime})>\varepsilon\right)\ll\|R_{h}f-f^{\prime}\|_{2}\leq\delta.$
@@
3. Proof of Theorem 1.5
In this section we show that, provided Condition ($\tilde{C}_{i}$) holds for some $1\leq i\leq d$, the operators $R_{k}$ enjoy the "strong sweeping out" property. The proof is similar to that of [JONES1991127][Theorem 3.2] and is a direct application of [BJR90, Theorem 3], which we recall below for the convenience of the reader.
@@
When $T_{k}=R_{k}$, the first two assumptions in Theorem 3.1 are trivially true, while the third one follows by our hypothesis in Theorem 1.5. Hence, it is enough for us to verify that ($\star$3.1) holds. Note also that, when $f$ is the characteristic function of a set in $X$, $M_{n}f$ coincides with $N_{\mathcal{B}^{\prime}}f$, where $\mathcal{B}^{\prime}=\{B_{k}\in\mathcal{B}:k\geq n\}$.
We first need a Rokhlin tower-type construction for aperiodic flows proved in [Lind], which we recall below.
@@
[Lind, Theorem 1] Let $U_{\boldsymbol{t}}=U_{t_{1},\ldots,t_{d}}$ be a $d$-dimensional measure-preserving aperiodic flow on $(X,\mu)$. Let $L_{1},\ldots,L_{d},\delta>0$ and let $Q=Q_{L}:=[0,L_{1})\times\ldots\times[0,L_{d})$. Then there exists a set $B\subset X$ with the following properties: